Solve by substitution.
3y = x + 5
3x - 9y = -10
3y = x + 5
3x - 9y = -10
=> -9y = -10 -3x
=> -9(x + 5) = -10 - 3x
=> -9x - 45 = -10 - 3x
=> -45 + 10 = -3x + 9x
=> -35 = 6x
=> -35/6 = x
3y = x + 5
=> 3y = -35/6 + 5
=> 3y = -35/6 +30/6
=> 3y = -5/6
=> y = -15/6

Answer by dead_famous1991(7)


3y = x + 5
3x - 9y = -10
3y = x + 5
3x - 9y = -10
=> -9y = -10 -3x
=> -9(x + 5) = -10 - 3x
=> -9x - 45 = -10 - 3x
=> -45 + 10 = -3x + 9x
=> -35 = 6x
=> -35/6 = x
3y = x + 5
=> 3y = -35/6 + 5
=> 3y = -35/6 +30/6
=> 3y = -5/6
=> y = -15/6